Consider a simple economy that produces two goods: pens and erasers. The following table shows the prices and quantities of the goods over a three-year period.

Year

Pens

Erasers

Price

Quantity

Price

Quantity

(Dollars per pen)

(Number of pens)

(Dollars per eraser)

(Number of erasers)

2016 1 110 2 150
2017 2 155 4 215
2018 3 120 4 180

Use the information from the preceding table to fill in the following table.

Year

Nominal GDP

Real GDP

GDP Deflator

(Dollars)

(Base year 2016, dollars)

2016
2017
2018

From 2017 to 2018, nominal GDP   , and real GDP   .

The inflation rate in 2018 was   .

Why is real GDP a more accurate measure of an economy's production than nominal GDP?

Nominal GDP is adjusted for the effects of inflation or deflation, whereas real GDP is not.

Real GDP measures the value of the goods and services an economy produces, but nominal GDP measures the value of the goods and services an economy consumes.

Real GDP is not influenced by price changes, but nominal GDP is.
Answer #1

GDP = Price * Quantity

Nominal GDP

2016 = 1*110 + 2*150 = 110 + 300 = $410

2017 = 2 * 155 + 4 * 215 = $1170

2018 = 3 * 120 + 4 * 180 = $1080

Real GDP

Base Year 2016

2016 = $410

2017 = 1 * 155 + 2 * 215 = $585

2018 = 1 * 120 + 2 * 180 = $480

GDP Deflator

(Nominal GDP / Real GDP ) * 100

2017 = (1170 / 585 ) * 100 = 200

2018 = ( 1080 / 480 ) * 100 = 225

Inflation Rate 2018 as compared to 2017 = (225 - 200) / 200 = 12.5%

GDP or Gross Domestic Product is the total value of all goods and services a country produces in a calendar year. Unlike nominal GDP, real GDP accounts for changes in price levels and provides a more accurate figure of economic growth.Real GDP is not influenced by price changes, but nominal GDP is. Real GDP is a measure of a country's total economic output, adjusted for price changes. Real GDP makes comparing GDP from year to year and from different years more meaningful because it shows comparisons for both the quantity and value of goods and services.